# The Tidal Torque Theory Revisited. I. Protohalo Angular Momentum

### v2

@@ Abstract
According to the Tidal Torque Theory (TTT), the angular momentum of haloes arises from the tidal torque produced on protohaloes by the mass fluctuations around them. That approach, initially developed assuming protohaloes as random overdense regions in the linear density field, was extended to the more realistic scenario that protohaloes are collapsing patches around peaks in the Gaussian-smoothed linear density field. But that extension faced two fundamental issues: 1) the unknown mass of collapsing patches marked by Gaussian peaks, and 2) the unknown ellipsoidal collapse time of those triaxial patches. Furthermore, the TTT strictly holds in linear regime only. This Paper is the first of a series of two devoted to revisiting the TTT and accurately calculating the halo angular momentum. In the present Paper we use the CUSP formalism fixing all those issues to deal with the TTT from a full peak model viewpoint, i.e. not only is the protohalo suffering the tidal torque identified to a peak, but the main mass fluctuation causing the tidal torque is also seen as a peak or a hole. This way we obtain a simple analytic expression for the Lagrangian protohalo AM, which can be readily implemented in galaxy formation models and be compared to the results of simulations.
@@ 1 INTRODUCTION
It is commonly accepted that haloes acquire their angular momentum (AM) as a consequence of the tidal torque caused by neighbouring mass fluctuations as suggested by Hoyle et al. (1949).
@@
Peebles (1969) applied this Tidal Torque Theory (TTT) to estimate the growth of the AM in spherical protohaloes (see also Thuan & Gott 1977). Doroshkevich (1970) refined the theory taking into account that protohaloes are triaxial and White (1984) completed this picture by taking into account that the tidal torque on triaxial protohaloes is due to the also non-spherical gravitational potential of the neighbouring mass distribution and that the principal axes of the inertia tensor of the protohalo and those of the spatial distribution of the surrounding matter (and of the tidal tensor it yields) are not aligned in general.
@@
Taylor expanding to second order the peculiar gravitational potential $\Phi({\bf x})$ at the centre of mass (c.o.m.) of the protohalo and integrating over the protohalo volume, White (1984) found that, in linear regime, the AM grows, keeping the spin direction fixed, according to
$J_{i}(t)\approx-a^{2}(t)\dot{D}(t)\,\epsilon_{ijk}{\bf T}_{jl}{\bf I}_{lk}$ (1)
@@
where $J_{i}(t)$ is the $i$th Cartesian component of the angular momentum J of the protohalo at the cosmic time $t$, $\epsilon_{ijk}$ is the fully antisymmetric Levi-Civita rank-three tensor, ${\bf T}_{ij}\equiv\partial^{2}\Phi/\partial x_{i}\partial x_{j}$ is the tidal (or deformation) tensor, i.e. the Hessian of $\Phi$, at the protohalo center, and ${\bf I}$ is the inertia tensor of the protohalo with respect to its c.o.m., both appropriately scaled to be independent of the arbitrary time when they are calculated. The growth of $J_{i}$ is then encoded in the factor $a^{2}(t)\dot{D}(t)$, where $a(t)$ is the cosmic scale factor and $D(t)$ is the linear growth factor of density perturbations, which in the Einstein-de Sitter (EdS) cosmology is proportional to $t$. This factor arises from the time dependence of the integral of the cross product of ${\bf r}\times\dot{\bf r}$ over all mass particles in the protohalo, taking into account that in linear regime the space stretches proportionally to $a(t)$ and the peculiar velocity of particles, $\dot{\bf r}$, is proportional to their peculiar acceleration, which factorises as $\dot{D}\nabla\Phi$. The remaining factor, which is thus time-independent, is the Lagrangian AM of the protohalo. Indeed, equation (1) accounts for the fact that, in linear regime, the spatial distribution of matter is kept unchanged except for the scaling factor $a(t)$ and that overdensities grow as $D(t)$ (Sugerman, Summers & Kamionkowski, 2000; Porciani, Dekel, & Hoffman, 2002a, b). In fact, it also approximately holds in mildly non-linear regime (Catelan & Theuns, 1996b).
@@
In the original Peebles’s form of the TTT, protohaloes were assumed to be simple overdense patches in the linear Gaussian random density field suffering top-hat spherical collapse, like in the Press-Schechter (1974; PS) and the ‘excursion set’ (ES; Bond et al. 1991) models. But in more realistic Doroshkevich’s and White’s picture, protohaloes are triaxial patches around local maxima (peaks) in the Gaussian-smoothed density field, which suffer ellipsoidal collapse. This motivated Heavens & Peacock (1988), Hoffman (1988) and Catelan & Theuns (1996a, b) to implement the TTT in the peak model. But that implementation faced two fundamental difficulties: the unknown extension (or mass) and collapse time of patches marked by triaxial Gaussian peaks. This forced those authors to adopt a few uncertain assumptions. In addition, their predictions are in the numerical form, difficult to implement in analytic models or to be compared with the results of simulations. In particular, (Hoffman, 1988) calculated the protohalo AM by means of Monte Carlo simulations, while Heavens & Peacock (1988) and Catelan & Theuns (1996a) derived it analytically, but could not avoid the final complex result to be numerical. Last but not least, the translation of the predicted protohalo AM to the final halo AM is affected by the general limitation that equation (1) does not hold in the final accelerated collapse and the virialisation (through shell-crossing) of the halo.
But the CUSP (ConflUent System of Peak trajectories) formalism introduced in Manrique & Salvador-Solé (1995, 1996) and Manrique et al. (1998) allows one to determine the extension and collapse time of triaxial ellipsoidal collapse, so it is time to revisit the TTT theory taking into account these improvements. Moreover, CUSP accurately reproduces the properties of (collapsed and virialised) haloes found in simulations Salvador-Solé & Manrique (2021), so it should hopefully also allow one to accurately calculate their AM. That is the aim of the present suite of two Papers.
In the present one, we concentrate in the calculation of the protohalo AM. Our approach is not only more accurate than previous ones regarding the above mentioned issues, but it also completes the peak viewpoint in the sense that not only is the protohalo identified to a peak but also the main source of the tidal torque is treated as such. As an extra bonus, we find a fully analytic expression of the Lagrangian AM of protohaloes, which can be readily used in the analytic modelling of structure formation and in the comparison with the results of simulations. In Paper II we use these results to derive accurate expressions for the AM and spin of final haloes as well as their specific angular momentum profile.
@@
The layout of the Paper is as follows. In Section 2 we remind the properties of protohaloes in the CUSP formalism. The properties of the neighbouring mass excess or defect responsible for the tidal torque acting on a protohalo are derived in Section 3. In Section 4 we use the previuous results to obtain the Lagrangian (time-invariant) AM of protohaloes which determines the growth of the Eulerian counterpart in linear and mildly non-linear regimes. Our results are discussed and summarised in Section 5.
@@ 2 Protohalo
In the peak model, protohaloes are collapsing patches traced by peaks (secondary maxima) in the linear random Gaussian density field. These collapsing patches are triaxial rather than spherical, so they undergo ellipsoidal collapse. The collapse time (along all three axes) $t$ depends in this case not only on the mass (or scale $R$) and density contrast $\delta$ of the peak, like in spherical collapse, but also on the curvature $x$ (or sharpness) of the peak, i.e. $-\nabla^{2}\delta({\bf x})$ scaled to the rms value of that quantity) as well as on its shape, i.e. ellipticity $e$ and prolateness $p$, $t=t(\delta,x,e,p,R)$. Nonetheless, as we will see shortly, the PDF of $e$, $p$ and, particularly, $x$ of peaks with density contrast $\delta$ at scale $R$ in the density field at $t_{\rm i}$ smoothed with a Gaussian window are sharply peaked, so we can assume them with fixed values at the respective maxima, $e_{\rm max}$, $p_{\rm max}$, $x_{\rm max}$. Therefore, given any halo mass definition (see below) fixing the mass $M$ of haloes at $t$, the corresponding ellipsoidal patches at $t_{\rm i}$ traced by triaxial peaks with a positive $\delta$ at the suited scales $R(M,t)$ will collapse essentially at the same time $t=t[\delta,x_{\rm max},e_{\rm max},p_{\rm max},R(M,t)]$, with $t$ a monotonic decreasing function of $\delta$ like in top-hat spherical collapse. In other words, in ellipsoidal collapse, there is indeed a one-to-one correspondence between haloes with $M$ at $t$ and peaks with $\delta$ and $R$ in the density field at $t_{\rm i}$ filtered with a Gaussian window.
In practice, the relations between $\delta$ and $R$ of peaks and $M$ and $t$ of haloes can be obtained by enforcing the consistency conditions that all the DM in the Universe at any $t$ must be locked inside haloes of different masses and that the mass $M$ of a halo must be equal to the volume-integral of its density profile (Juan et al., 2014). Specifically, if we write the density contrast $\delta$ and scale $R$ of peaks at $t_{\rm i}$ in the general form
@@
where $\delta_{\rm c}^{\rm th}(t)$ is the critical linearly extrapolated density contrast for top-hat spherical collapse at $t$ equal to $3(12\pi)^{2/3}/20\approx 1.686$ in the EdS cosmology, $D(t)$ is the linear growth factor, equal to the cosmic scale factor $a(t)$ in that universe, and $\sigma_{i}(R,t,t_{\rm i})$ is the Gaussian $i$th-order spectral moment on scale $R$, then the functions $r_{\delta}$ and $r_{\sigma}$ are well-fitted, in all cases analysed, by the analytic expressions (see Salvador-Solé & Manrique 2024)
@@
where $\nu^{\rm th}(M,t)\equiv\delta^{\rm th}(t,t_{\rm i})/\sigma^{\rm th}_{0}(M,t_{\rm i})=\delta^{\rm th}_{\rm c}(t)/\sigma^{\rm th}_{0}(M,t)$ is the peak height in top-hat smoothing. The values of the coefficients $d$, $s_{0}$, $s_{1}$, $s_{2}$ and $A$ for some relevant cosmologies (Table 2) and halo mass definitions are given in Table 1. The previous relations imply in turn the following $t_{\rm i}$-independent relation between the Gaussian and top-hat peak heights is
@@
Equation (3) defines the Gaussian scale $R(M,t,t_{\rm i})$ that corresponds to the mass $M$ of the halo at $t$ in terms of the top-hat 0th-order spectral moment at $t$, $\sigma_{0}^{\rm th}(M,t)$ and it shows that the Gaussian scale $R$ is not simply proportional to the top-hat one $R^{\rm th}$ satisfying $M=4\pi/3(R^{\rm th})^{3}\bar{\rho}(t_{\rm i})$. But, in the case of power-law power spectra, $P(k)=P_{0}k^{n}$, there is a more direct relation between the Gaussian scale $R$ and of the top-hat one $R^{\rm th}$ delimiting physically the protohalo of mass $M$. Indeed, the definition of the $j$th-order spectral moment for Gaussian and top-hat filters, denoted with index $f$ equal to ‘blank’ and to ‘th’, respectively,
$(\sigma_{j}^{f})^{2}(R^{f})=\frac{P_{0}}{2\pi^{2}(R^{f})^{n+3+2{\rm j}}}\int_{0}^{\infty}{\rm d}x\,x^{n+2(1+{\rm j})}\,(W^{f})^{2}(x),$ (7)
@@
where $Q^{2}_{0}\equiv\int_{0}^{\infty}{\rm d}x\,x^{n+2}\,W^{2}(x)/\int_{0}^{\infty}{\rm d}x\,x^{n+2}\,(W^{\rm th})^{2}(x)$, and $W$ and $W^{\rm th}$ are the Fourier transforms of the Gaussian and top-hat filters, respectively. Thus, taking into account that the CDM spectrum in the galaxy mass range is locally of the power-law form (with $n\approx-1.75$ and $Q_{0}\approx 0.6$ in the galaxy mass range), we can also adopt the relations (8)-(9) as a useful approximation in the case of the CDM spectrum. It is worth mentioning that, in the case of virial masses ($M_{\rm vir}$), $r_{\sigma}$ is a function of $M$ alone, $r_{\sigma}(M)=1+c/\sigma_{0}^{\rm th}(M)$, where $\sigma_{0}^{\rm th}(M)$ is the top-hat 0th order spectral moment at the present time and $c=0.14$ and 0.10 in the WMAP7 and Planck14 cosmologies, respectively (Salvador-Solé & Manrique, 2024), so $R$ is a function of $M$ alone, of the order of $R\approx 2R^{\rm th}$ in the galactic halo mass range.
@@
The average number density of peaks at $t_{\rm i}$ with $\nu$, $x$, $e$ and $p$ and oriented (relative at the present stage to some arbitrary reference) according to the Euler angles $\alpha$, $\beta$ and $\kappa$ at the scale $R$ per infinitesimal range of all these arguments is (BBKS)
@@
where $\gamma\equiv\sigma_{1}^{2}/(\sigma_{0}\sigma_{2})$ and $R_{\star}\equiv\sqrt{3}\,\sigma_{1}/\sigma_{2}$ ($\gamma=[(n+3)/(n+5)]^{1/2}$ and $R_{\star}=[(n+5)/6]^{-1/2}R$ for power-law spectra),
@@
is the joint probability function of ellipticities, $e=(\lambda_{1}-2\lambda_{2}+\lambda_{3})/(2x)$ ($e\geq 0$), and prolatenesses $p=(\lambda_{1}-\lambda_{3})/(2x)$ ($-e\leq p\leq e$), where $\lambda_{i}\geq\lambda_{2}\geq\lambda_{3}$ are the eigenvalues, with positive sign, of the second derivative tensor ($\lambda_{1}+\lambda_{2}+\lambda_{3}=x$, so there are 2 degrees of freedom only), which, for large $x$ as corresponding to galactic haloes, is very nearly Gaussian with corresponding means $\bar{e}=1/\{\sqrt{5}x[1+6/(5x^{2})]^{1/2}\}$ and $\bar{p}=6/\{5x^{4}[1+6/(5x^{2})]^{2}\}$ and dispersions $\sigma_{\rm e}=\bar{e}/\sqrt{6}$ and $\sigma_{\rm p}=\bar{e}/\sqrt{3}$.
As mentioned, the $x$-PDF given by the second factor on the right of equation (10) is very sharply peaked and all peaks with $\nu$ at $R$ have nearly the same curvature, [^1] The $e$- and $p$-PDFs are also sharply peaked, but the protohalo AM strongly depends on the triaxial shape and orientation of the corresponding peaks, so we will not to make a similar approximation for these quantities. equal to its maximum or mean value
@@
In addition, the average of any function $f(x)$ of the peak curvature $x$ is very nearly equal to $f(\langle x\rangle)$, which can be used in the above mentioned mean peak ellipticity and prolateness and the corresponding dispersions.
Thus, we can drop the curvature from the list of compelling peak properties and deal with the average number density of peaks at $t_{\rm i}$ with $C=\{e,p,\alpha,\beta,\kappa\}$ at $R$, given by the integral over $x$ of the number density (10),
@@
As mentioned, collapsing patches are ellipsoidal, implying that peaks are triaxial. In fact, $e$ and $p$ are the ellipticity and prolateness of peaks, not of protohaloes. The physical (top-hat) semi-axes of the protohalo, $a_{1}$, $a_{2}$ and $a_{3}$ satisfy the condition $a_{1}a_{2}a_{3}=(R^{\rm th})^{3}$ (there are also two degrees of freedom in the $a_{i}$ values) and are related to the eigenvalues of the peak through
@@
where $\hat{a}_{i}$ are the scaled semi-axes defined as
@@
Consequently, the Lagrangian (time-independent) inertia tensor I of the ellipsoidal protohalo with uniform density (equal to $\bar{\rho}(t_{\rm i})$ at $t_{\rm i}$) relative to the c.o.m. in Cartesian coordinates oriented along the principal axes is
${\bf I}=\frac{Ma^{2}(t)}{5\,a^{2}(t_{\rm i})}\left(\begin{matrix}\iota_{1}&0&0\\ 0&\iota_{2}&0\\ 0&0&\iota_{3}\end{matrix}\right),$ (20)
where $\iota_{1}\equiv a_{2}^{2}+a_{3}^{2}$, $\iota_{2}\equiv a_{1}^{2}+a_{3}^{2}$ and $\iota_{3}\equiv a_{2}^{2}+a_{1}^{2}$. Note that by dividing the inertia tensor by the $a^{2}(t_{\rm i})$, the tensor I does not depend on $t_{\rm i}$ as desired.

### v5

@@ Abstract
According to the Tidal Torque Theory (TTT), the angular momentum (AM) of dark matter haloes arises from the tidal torque suffered by ellipsoidal protohaloes due to their surrounding mass distribution. Unfortunately, the practical implementation of the TTT in the peak model of structure formation, where protohaloes are, indeed, ellipsoidal patches around triaxial peaks (maxima) in the Gaussian-smoothed density field, faces some fundamental issues. First, the mass and collapse time of patches marked by Gaussian peaks are poorly determined. Second, the growth of the resulting AM is only known in linear regime. This Paper is the first of two devoted to revisiting the TTT and accurately calculating the rotational halo properties it predicts. Here we use the CUSP formalism fixing the above mentioned indeterminacy in the protohalo AM. The peak viewpoint is fully accomplished in the sense that not only is the protohalo identified to a peak, but also the mass fluctuations causing the tidal torque are identified to peaks or holes (local minima). This way we obtain a very simple analytic expression for the mean (and median) Lagrangian protohalo AM, which can be readily implemented in galaxy formation models and be compared to the results of simulations.
@@ 1 INTRODUCTION
It is commonly accepted that dark matter (DM) haloes acquire their angular momentum (AM) as a consequence of the tidal torque caused by neighbouring mass fluctuations, as suggested by Hoyle et al. (1949).
@@
Peebles (1969) applied this Tidal Torque Theory (TTT) to estimate the growth of the AM in spherical protohaloes (see also Thuan & Gott 1977). Doroshkevich (1970) refined the theory taking into account that protohaloes are ellipsooidal and White (1984) completed this picture by taking into account that the tidal torque is due to the also non-spherical gravitational potential of the neighbouring mass distribution, being the principal axes of the inertia and tidal tensors not aligned in general.
@@
Taylor expanding to second order the peculiar gravitational potential $\Phi({\bf x})$ at the centre of mass (c.o.m.) of the protohalo and integrating over its volume, White (1984) found that, in linear regime, the AM grows, keeping the spin orientation fixed, according to
$J_{i}(t)\approx-a^{2}(t)\dot{D}(t)\,\epsilon_{ijk}{\bf T}_{jl}{\bf I}_{lk},$ (1)
@@
where $J_{i}(t)$ is the $i$th Cartesian component of the angular momentum J of the protohalo at the cosmic time $t$, $\epsilon_{ijk}$ is the fully antisymmetric Levi-Civita rank-three tensor, ${\bf T}_{ij}\equiv\partial^{2}\Phi/\partial x_{i}\partial x_{j}$ is the tidal (or deformation) tensor, i.e. the Hessian of the potential $\Phi$, at the protohalo c.o.m. and ${\bf I}$ is the inertia tensor of the protohalo with respect to that point, with both quantities appropriately scaled so as to be independent of the arbitrary time when they are calculated. The growth of $J_{i}$ is thus encoded in the factor $a^{2}(t)\dot{D}(t)$, where $a(t)$ and $D(t)$ are the cosmic scale and linear growth factor, respectively. This way, $J_{i}$ grows as $t$ in the Einstein-de Sitter (EdS) cosmology.
Factor $a^{2}(t)\dot{D}(t)$ arises from the time dependence of the integral of the cross product of ${\bf r}\times\dot{\bf r}$ over the protohalo mass particles, taking into account that in linear regime the space stretches proportionally to $a(t)$ and the peculiar particle velocity $\dot{\bf r}$ is proportional to their peculiar acceleration, which factorises as $\dot{D}\nabla\Phi$. The remaining time-independent quantity is the Lagrangian AM of the linear protohalo.
Strictly speaking, equation (1) holds for protohaloes in linear regime only. However, it also approximately holds in mildly non-linear regime (Catelan & Theuns, 1996b). In addition, simulations show that the AM of relaxed haloes is roughly consistent with such a growth until the final accelerated collapse of the protohalo (e.g. Sugerman, Summers & Kamionkowski 2000; Porciani, Dekel, & Hoffman 2002a, b). But one cannot be more precise because the translation from the protohalo AM to the final AM of relaxed haloes involves highly non-linear effects, such as the accelerated collapse and virialisation (through shell-crossing) of the objects and the presence of major mergers, that are very hard to properly address.
@@
Moreover, even focusing on protohaloes in linear regime, the TTT is not a closed subject. In the PS (Press & Schechter, 1974) and ‘excursion set’ (ES; Bond et al. 1991) models of structure formation where protohaloes are overdense patches undergoing top-hat spherical collapse, the protohalo AM does not grow. It only does provided protohaloes are ellipsoidal, as in the more realistic peak model introduced by Doroshkevich and White. This motivated Heavens & Peacock (1988), Hoffman (1988) and Catelan & Theuns (1996a, b) to implement the TTT in such a model where protohaloes correspond to triaxial density maxima in the Gaussian-smoothed random Gaussian density field. Unfortunately, this implementation faced the fundamental issue that the exact extension (or mass) and collapse time of the collapsing patches encompassed by Gaussian peaks are unknown, which forced those authors to adopt some uncertain assumptions. In addition, calculations were quite complex and the final results could only be in numerical form. In particular, Hoffman (1988) used Monte Carlo simulations, while the derivations by Heavens & Peacock (1988) and Catelan & Theuns (1996a) are analytical, though involving numerical integrations. Consequently, their predictions are difficult to compare with the results of simulations and to implement in analytic models.
But all these shortcomings can in principle be overcome thanks to the ConflUent System of Peak trajectories (CUSP) formalism introduced in Manrique & Salvador-Solé (1995, 1996) and Manrique et al. (1998). This formalism allows one to determine the mass and collapse time of patches marked by Gaussian peaks and to deal with the effects of collapse and virialisation and of major mergers as well (Salvador-Solé & Manrique, 2021). Thus, it should hopefully also enable one to properly address the halo AM in the TTT. This is precisely the aim of the series of two Papers started here.
In the present Paper, we concentrate in revisiting the TTT and accurately calculate the AM of linear protohaloes in the peak model. Making use of the CUSP formalism, we solve the above mentioned fundamental issues of this model. In addition, the problem is dealt with from a full peak viewpoint in the sense that, not only is the protohalo identified to a peak but also the main source of the tidal torque acting on it is treated as a peak or a hole (local density minima). As a bonus, we find a very practical fully analytic expression of the Lagrangian protohalo AM. These results are used in Paper II to accurately derive all rotational halo properties.
@@
The layout of the Paper is as follows. In Section 2 we remind the protohalo properties in the CUSP formalism. The properties of the neighbouring mass excess or default responsible for the tidal torque acting on a protohalo are derived in Section 3. In Section 4 we use the previous results to obtain the (time-invariant) Lagrangian protohalo AM which determines the growth of the Eulerian counterpart in linear and mildly non-linear regimes. Our results are discussed and summarised in Section 5.
@@ 2 Protohalo
In the peak model, protohaloes are collapsing patches traced by peaks in the linear random Gaussian density field. These collapsing patches are triaxial (like their associated peaks) rather than spherical, so they undergo ellipsoidal collapse. Their collapse time (along all three axes) $t$ then depends not only on the mass (or scale $R$) and density contrast $\delta$ of the peak, like in spherical collapse, but also on its curvature (or sharpness) $x$, defined as $-\nabla^{2}\delta({\bf x})$ scaled to the rms value of that quantity, as well as on its shape, i.e. ellipticity $e$ and prolateness $p$, i.e. $t=t(\delta,x,e,p,R)$. Nonetheless, as we will see shortly, the PDF of $e$, $p$ and, particularly, $x$ of peaks with density contrast $\delta$ at scale $R$ in the Gaussian-smoothed density field at $t_{\rm i}$ are sharply peaked, so we can assume them with fixed values at the respective maxima, $e_{\rm max}$, $p_{\rm max}$, $x_{\rm max}$. Therefore, given any halo mass definition (see below) fixing the mass $M$ of haloes at $t$, the corresponding ellipsoidal patches at $t_{\rm i}$ traced by triaxial peaks with a positive $\delta$ at the suited scales $R(M,t)$ will collapse essentially at the same time, with $t$ a monotonic decreasing function of $\delta$ like in top-hat spherical collapse. In other words, in Gaussian ellipsoidal collapse like in top-hat spherical one, there is a correspondence between haloes with $M$ at $t$ and peaks with $\delta$ and $R$ in the Gaussian-smoothed density field at $t_{\rm i}$. [^1] This correspondence is not one-to-one because some peaks lie within others with identical density contrast at a higher scale, so they are swallowed by their host (and become subhaloes) before accomplishing full collapse. However, that is irrelevant for our purposes here.
The relations between $\delta$ and $R$ of peaks and $M$ and $t$ of the corresponding haloes can be obtained by enforcing the consistency conditions that all the DM in the Universe at any time $t$ is locked in haloes of different masses and that the mass $M$ of a halo is equal to the volume-integral of its density profile times the squared radius (Juan et al., 2014). Specifically, if we write the density contrast $\delta$ and scale $R$ of peaks at $t_{\rm i}$ in the general form
@@
where $\delta_{\rm c}^{\rm th}(t)$ is the critical linearly extrapolated density contrast for top-hat spherical collapse at $t$, equal to $3(12\pi)^{2/3}/20\approx 1.686$ in the EdS cosmology, $D(t)$ is the linear growth factor, equal to the cosmic scale factor $a(t)$ in that universe, and $\sigma_{i}(R,t,t_{\rm i})$ is the Gaussian $i$th-order spectral moment on scale $R$, then the functions $r_{\delta}$ and $r_{\sigma}$ are well-fitted, in all cases analysed, by the analytic expressions (see Salvador-Solé & Manrique 2024)
@@
where $\nu^{\rm th}(M,t)\equiv\delta^{\rm th}(t,t_{\rm i})/\sigma^{\rm th}_{0}(M,t_{\rm i})=\delta^{\rm th}_{\rm c}(t)/\sigma^{\rm th}_{0}(M,t)$ is the peak height in top-hat smoothing. The values of the coefficients $d$, $s_{0}$, $s_{1}$, $s_{2}$ and $A$ for some relevant cosmologies (Table 2) and halo mass definitions are given in Table 1. The previous relations imply in turn the following $t_{\rm i}$-independent relation between the Gaussian and top-hat peak heights
@@
Equation (3) defines the Gaussian scale $R(M,t,t_{\rm i})$ that corresponds to the mass $M$ of haloes at $t$ in terms of the top-hat 0th-order spectral moment at $t$, $\sigma_{0}^{\rm th}(M,t)$ and it shows that the Gaussian scale $R$ is not simply proportional to the top-hat one $R^{\rm th}$ satisfying $M=4\pi/3(R^{\rm th})^{3}\bar{\rho}(t_{\rm i})$. However, in the case of power-law power spectra, $P(k,t)=P_{0}(t)k^{n}$, there is a more direct relation between the Gaussian scale $R$ and of the top-hat one $R^{\rm th}$ delimiting physically the protohalo of mass $M$. Indeed, the definition of the $j$th-order spectral moment for Gaussian and top-hat filters, denoted with index $f$ equal to ‘blank’ and to ‘th’, respectively,
$(\sigma_{j}^{f})^{2}(R^{f})=\frac{P_{0}D(t_{0})/D(t)}{2\pi^{2}(R^{f})^{n+3+2{\rm j}}}\int_{0}^{\infty}{\rm d}x\,x^{n+2(1+{\rm j})}\,(W^{f})^{2}(x),$ (7)
@@
where $Q^{2}_{0}\equiv\int_{0}^{\infty}{\rm d}x\,x^{n+2}\,W^{2}(x)/\int_{0}^{\infty}{\rm d}x\,x^{n+2}\,(W^{\rm th})^{2}(x)$, and $W$ and $W^{\rm th}$ are the Fourier transforms of the Gaussian and top-hat filters, respectively. Thus, taking into account that the CDM spectrum in the galaxy mass range is locally of the power-law form (with $n\approx-1.75$ and $Q_{0}\approx 0.6$ in the galaxy mass range), we can also adopt the relations (8)-(9) as a useful approximation in CDM. It is worth mentioning that, in the case of virial masses ($M_{\rm vir}$ defined in the caption of Table 1), $r_{\sigma}$ is a function of $M$ alone, $r_{\sigma}(M)=1+c/\sigma_{0}^{\rm th}(M)$, where $\sigma_{0}^{\rm th}(M)$ is the top-hat 0th order spectral moment at the present time and $c=0.14$ and 0.10 in the WMAP7 and Planck14 cosmologies, respectively (Salvador-Solé & Manrique, 2024). Thus, $R$ is, in this case, a function of $M$ alone of the order of $R\approx 2R^{\rm th}$.
In what fallows, we will be concerned with to different cosmic times apart from the arbitrary initial one $t_{\rm i}$: one to describe the evolution of the protohalo AM and the other one referring to the protohalo collapse time. To avoid any confusion in the notation, $t$ will be reserved to the former, while for the latter, which appears much less frequently, we will use $t_{\rm c}$.
@@
The average number density of peaks at $t_{\rm i}$ with $\nu$, $x$, $e$ and $p$ and oriented according to the Euler angles $\alpha$, $\beta$ and $\kappa$ at the scale $R$ per infinitesimal range of all these arguments is (BBKS)
@@
where $\gamma\equiv\sigma_{1}^{2}/(\sigma_{0}\sigma_{2})$ $R_{\star}\equiv\sqrt{3}\,\sigma_{1}/\sigma_{2}$ ($\gamma=[(n+3)/(n+5)]^{1/2}$ ($R_{\star}=[(n+5)/6]^{-1/2}R$ for power-law spectra),
@@
is the joint probability function of ellipticities, $e=(\lambda_{1}-2\lambda_{2}+\lambda_{3})/(2x)$ ($e\geq 0$), and prolatenesses $p=(\lambda_{1}-\lambda_{3})/(2x)$ ($-e\leq p\leq e$), where $\lambda_{i}\geq\lambda_{2}\geq\lambda_{3}$ are the eigenvalues, with positive sign, of the second order derivative tensor ($\lambda_{1}+\lambda_{2}+\lambda_{3}=x$, so there are 2 degrees of freedom only), which, for large $x$ as corresponding to galactic haloes, are nearly Gaussian with corresponding means $\bar{e}=1/\{\sqrt{5}x[1+6/(5x^{2})]^{1/2}\}$ and $\bar{p}=6/\{5x^{4}[1+6/(5x^{2})]^{2}\}$ and dispersions $\sigma_{\rm e}=\bar{e}/\sqrt{6}$ and $\sigma_{\rm p}=\bar{e}/\sqrt{3}$ (BBKS).
As mentioned, the $x$-PDF given by the second factor on the right of equation (10) is very sharply peaked and all peaks with $\nu$ at $R$ have nearly the same curvature, [^2] Even though the $e$- and $p$-PDFs are also quite sharply peaked, the protohalo AM crucially depends on the triaxial shape and orientation of the corresponding peaks. Thus, we cannot make such an approximation for these quantities. equal to its maximum or mean value
@@
In addition, the average of any function $f(x)$ of the peak curvature is very nearly equal to $f(\langle x\rangle)$, which can be used in the above mentioned mean peak ellipticity and prolateness and the corresponding dispersions.
Thus, we can marginalise the curvature and work with the average number density of peaks at $t_{\rm i}$ with $C=\{e,p,\alpha,\beta,\kappa\}$ at $R$,
@@
As mentioned, collapsing patches are ellipsoidal, but the previous quantities $e$ and $p$ are the ellipticity and prolateness of peaks, not of protohaloes. The physical (top-hat) semi-axes, $a_{1}$, $a_{2}$ and $a_{3}$ of the protohalo satisfy the condition $a_{1}a_{2}a_{3}=(R^{\rm th})^{3}$ (there are also two degrees of freedom in the $a_{i}$ values) and are related to the eigenvalues of the correspondng peak through
@@
where we have used the scaled semi-axes, $\hat{a}_{i}$, defined as
@@
Consequently, the Lagrangian inertia tensor I at $t$ of the ellipsoidal protohalo with uniform density $\bar{\rho}(t_{\rm i})$ relative to its c.o.m. in Cartesian coordinates oriented along the principal axes is
${\bf I}=\frac{M}{5\,a^{2}(t_{\rm i})}\left(\begin{matrix}\iota_{1}&0&0\\ 0&\iota_{2}&0\\ 0&0&\iota_{3}\end{matrix}\right),$ (20)
where $\iota_{1}\equiv a_{2}^{2}+a_{3}^{2}$, $\iota_{2}\equiv a_{1}^{2}+a_{3}^{2}$ and $\iota_{3}\equiv a_{2}^{2}+a_{1}^{2}$. We remark that, for simplicity, we have assumed that protohaloes are linear at all relevant scales, which implies that their density is everywhere essentially equal to $\bar{\rho}(t_{\rm i})$. Strictly speaking, protohaloes are not homogeneous, their typical density profile increases downwards until becoming non-linear at a small enough scale (Salvador-Solé et al. 2012a; Salvador-Solé & Manrique 2021), so their real inertia tensor may substantially deviate from that given in equation (20). This point will be properly addressed in Paper II.